OpenStudy (anonymous):

given a right triangle with an acute angle, if tangent angle = cotangent angle what would this triangle look like

OpenStudy (alexwee123):

a 45-45-90 triangle?

jimthompson5910 (jim_thompson5910):

tan(x) = opp/adj cot(x) = adj/opp So if tan(x) = cot(x), then opp/adj = adj/opp opp^2 = adj^2 opp = adj which means that this triangle has legs that are equal in length. So you would have a 45-45-90 degree triangle.

OpenStudy (anonymous):

 Looking at the drawing, \[\tan(x)=\frac{a}{b}\] \[\cot(x)=\frac{b}{a}\] since \[\tan(x)=\cot(x)\] \[\frac{a}{b}=\frac{b}{a}\] simplify, so \[a^2=b^2\] so \[a=b\] The legs of the triangle are the same length By numerous properties and definitiongs that you've already learned, since the legs of the right triangle are the same length, then the acute angles are the same, and this happens only in a 45-45-90 triangle